# Rational RG flow, extension, and Witt class

### v2

@@ Abstract
Consider a renormalization group flow preserving a pre-modular fusion category $\mathcal{S}_{1}$. If it flows to a rational conformal field theory, the surviving symmetry $\mathcal{S}_{1}$ flows to a pre-modular fusion category $\mathcal{S}_{2}$ with monoidal functor $F:\mathcal{S}_{1}\to\mathcal{S}_{2}$. By clarifying mathematical (especially category theoretical) meaning of renormalization group domain wall/interface or boundary condition, we find the hidden extended vertex operator (super)algebra gives a unique (up to braided equivalence) completely $(\mathcal{S}_{1}\boxtimes\mathcal{S}_{2})^{\prime}$-anisotropic representative of the Witt equivalence class $[\mathcal{S}_{1}\boxtimes\mathcal{S}_{2}]$. The mathematical conjecture is supported physically, and passes various tests in concrete examples including non/unitary minimal models, and Wess-Zumino-Witten models. In particular, the conjecture holds beyond diagonal cosets.
@@ 1 Introduction and summary
Conjecture. Let $\mathcal{S}_{\text{UV}}$ be the surviving pre-modular fusion category and $\mathcal{S}_{\text{IR}}$ be its image pre-modular fusion subcategory in the infrared diagonal rational conformal field theory. The (simple) rational renormalization group flow stands in bijection with a unique (up to braided equivalence) completely $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}$-anisotropic representative of the Witt equivalence class $[\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$
{Rational RG flow with $\{\text{Rational RG flow with }\mathcal{S}_{\text{UV}}\}\cong\{\text{Completely }(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}\text{-anisotropic representative of }[\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]\}.$ (1.1)
@@
Remark. If the surviving pre-MFC $\mathcal{S}_{\text{UV}}$ is non-degenerate, i.e., modular, the conjecture claims the representative of $[\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$ is completely anisotropic. Equivalently, the connected étale algebra $A\in\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}$ giving it is the maximal one.
Remark. The unique (up to equivalence) existence of completely $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}$-anisotropic representative of $[\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$ signals the universality. Since universality is usually defined in terms of Kan extension, it would be interesting to see the explicit relation between the conjecture and the axiom [14].
@@
Relatedly, when a rational RG flow admits RG domain wall/interface or boundary condition, the conjecture is trivially true by the bijection. Nontrivial statement of the conjecture is that all rational RG flows stand in bijection with completely $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}$-anisotropic representative of $[\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$.
@@
  • Second, the largest Frobenius-Perron dimension $\text{FPdim}_{\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}}(A)$ of a connected étale algebra $A$ is energetically favored. To see this, we first recall the largest Frobenius-Perron dimension minimizes the Frobenius-Perron dimension of the category $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})_{A}^{0}$ of dyslectic (right) $A$-modules. More precisely, in order to use the formula (A.7), we first remove the transparent simple objects. The resulting pre-MFC $\mathcal{B}$ no longer have nontrivial transparent simple objects (mathematically, this is denoted $A\cap(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}=1$), and it is modular. Since the connected étale algebra does not contain nontrivial transparent simple object, $A$ remains intact. For the MFC $\mathcal{B}$, we can apply the formula to get the smallest Frobenius-Perron dimension $\text{FPdim}(\mathcal{B})$. In pseudo-unitary fusion categories, the Frobenius-Perron dimension equals categorical dimension $\text{FPdim}(\mathcal{B})=D(\mathcal{B})^{2}$. An MFC with categorical dimension $D$ is known to contribute entropy [38, 39] by $-\ln D$. At temperature $T$ (defined as the length $1/T$ of the compactified Euclidean time), an MFC with categorical dimension $D$ thus contributes free energy by
@@
[…]
@@ 2.1 Unitary discrete series
[…]
@@
Indeed, in view of anyon condensation [45, 46, 47], one finds $A\in\mathcal{S}$ can be condensed. [^5] The simple objects of $\mathcal{S}_{A}$ are given by $m_{1}\cong A,\quad m_{2}\cong x\oplus X\oplus wW,\quad$ $m_{3}\cong yY\oplus zZ,\quad m_{4}\cong yZ\oplus zY,$ $m_{5}\cong y\oplus zX\oplus yW\oplus zW,\quad m_{6}\cong Y\oplus xZ\oplus wY\oplus wZ,\quad$ $m_{7}\cong z\oplus yX\oplus yW\oplus zW,\quad m_{8}\cong Z\oplus xY\oplus wY\oplus wZ,$ $m_{9}\cong w\oplus W\oplus wX\oplus xW\oplus wW,\quad$ $m_{10}\cong yY\oplus zZ\oplus yZ\oplus zY.$ The first four simple objects are deconfined, and the other six are confined. The deconfined simple objects have conformal dimensions $00,0,\frac{1}{2},$ 0 mod 1. We also found a non-negative integer matrix representation (NIM-rep), but we omit them. The result tells us
@@
$[(\mathcal{S}_{A}^{0})_{A^{\prime}}^{0}]=[\mathcal{S}_{A}^{0}]=[\mathcal{S}].$
Therefore, $[(\mathcal{S}_{A}^{0})_{A^{\prime}}^{0}]$ gives a unique (up to braided equivalence) completely anisotropic representative of the Witt equivalence class $[\mathcal{S}]$. This example further supports our conjecture. [^7] In order to make this example mathematically rigorous, we still need to prove $A\cong 1\oplus xX\oplus wW$ is a (connected) étale algebra. Since $F$-symbols of $su(2)_{4}$ seem unknown, we could not prove this. However, mathematicians would be able to prove this possibly with some codes.
@@ 2.3 Wess-Zumino-Witten model
where $g$ is the dual Coxeter number. The primary operators (and hence symmetry objects) are labeled by affine weights ${\hat{\mu}}$ in the affine dominant weights
@@ 2.4 $E$-type minimal model $(A_{10},E_{6})$
The $\mathbb{Z}/2\mathbb{Z}$ simple object in Ising always has conformal dimension $\frac{1}{2}$ mod 1. Thus, it trivially satisfies the half-integer condition. Let us look at the non-invertible simple object. It can have conformal dimensions $\frac{2n+1}{16}$ with $n=0,1,\dots,7$. In UV, it has $h_{y}=\frac{31}{16}$. For its image $Y=F(y)$ to satisfy the half-integer condition, the IR conformal dimension should be either $\frac{1}{16}$ or $\frac{9}{16}$. They give MFCs with additive central charges $\frac{1}{2},-\frac{7}{2}$ mod 8, respectively. For the latter to be unitary, the smallest additive central charge is $\frac{9}{2}$, which is larger than UV. This rules out the latter possibility as follows. All CFTs with central charge $10<c<$ 1 have been classified [24], and they are minimal models. None of minimal models has Ising with $h=\frac{9}{16}$ for the non-invertible simple object, and central charge smaller than $c_{\text{UV}}=\frac{21}{22}$. Therefore, the only possibility is Ising with $h=\frac{1}{16}$ for the non-invertible simple object. The only possibility is the critical Ising model $M(4,3)$.
@@ A.1 Category
:=$\mathcal{B}^{\prime}:=\{b_{i}\in\mathcal{B}|\forall b_{j}\in\mathcal{B},\ c_{b_{j},b_{i}}\circ c_{b_{i},b_{j}}\cong id_{b_{i}\otimes b_{j}}\}.$ (A.4)
A BFC is called symmetric if all simple objects are transparent. On the other extreme, a BFC is called non-degenerate if the identity object is the only transparent object, =$\mathcal{B}^{\prime}=\{1\}$. A non-degenerate BFC $\mathcal{B}\neq\text{Vect}_{\mathbb{C}}$ is called prime if there is no proper non-degenerate braided fusion subcategory other than $\text{Vect}_{\mathbb{C}}$. A generic non-degenerate BFC $\mathcal{B}\neq\text{Vect}_{\mathbb{C}}$ is a Deligne tensor product of prime non-degenerate braided fusion subcategories [26]
@@ A.3 Algebra
Let $\mathcal{B}_{1,2}$ be non-degenerate BFCs. They are called Witt equivalent if there exist fusion categories $\mathcal{C}_{1,2}$ such that $\mathcal{B}_{1}\boxtimes\mathcal{Z}(\mathcal{C}_{1})\simeq\mathcal{B}_{2}\boxtimes\mathcal{Z}(\mathcal{C}_{2})$ is a braided equivalence [26]. This introduces equivalence classes to non-degenerate BFCs. The Witt equivalence class of $\mathcal{B}$ is denoted by $[\mathcal{B}]$. For a non-degenerate BFC $\mathcal{B}$ and a connected étale algebra $A\in\mathcal{B}$, we have [26]
$[\mathcal{B}]=[\mathcal{B}_{A}^{0}].$ (A.8)
Let $\mathcal{B}_{1,2}$ be non-degenerate BFCs over a symmetric fusion category $\mathcal{E}$ such that the corresponding braided monoidal functors $\mathcal{E}\to\mathcal{B}_{i}\ (i=1,2)$ are fully faithful. The $\mathcal{B}_{1,2}$ are called Witt equivalent if there exist fusion categories $\mathcal{C}_{1,2}$ over $\mathcal{E}$ such that $\mathcal{B}_{1}\boxtimes_{\mathcal{E}}\mathcal{Z}(\mathcal{C}_{1},\mathcal{E})\simeq\mathcal{B}_{2}\boxtimes_{\mathcal{E}}\mathcal{Z}(\mathcal{C}_{2},\mathcal{E})$ is a braided equivalence. It introduces an equivalence relation to non-degenerate BFCs over $\mathcal{E}$. The Witt equivalence class of $\mathcal{B}$ is denoted $[\mathcal{B}]$. The set of Witt equivalence classes of non-degenerate BFCs over $\mathcal{E}$ is denoted by $\mathcal{W}(\mathcal{E})$. For a non-degenerate BFC $\mathcal{B}$ over $\mathcal{E}$ and a connected étale algebra such that $A\cap\mathcal{E}=1$, i.e., $A$ does not contain simple objects in $\mathcal{E}$ except the identity object, we have [27]
$[\mathcal{B}]=[\mathcal{B}_{A}^{0}].$ (A.9)

### v3

@@ Abstract
Consider a renormalization group flow preserving a pre-modular fusion category $\mathcal{S}_{1}$. If it flows to a rational conformal field theory, the surviving symmetry $\mathcal{S}_{1}$ flows to a pre-modular fusion category $\mathcal{S}_{2}$ with monoidal functor $F:\mathcal{S}_{1}\to\mathcal{S}_{2}$. By clarifying mathematical (especially category theoretical) meaning of renormalization group domain wall/interface or boundary condition, we find the hidden extended vertex operator (super)algebra gives a unique (up to braided equivalence) completely $(\mathcal{S}_{1}\boxtimes\mathcal{S}_{2})^{\prime}$-anisotropic representative of the Witt equivalence class $\mathcal{S}_{1}\boxtimes\mathcal{S}_{2}]$. The mathematical conjecture is supported physically, and passes various tests in concrete examples including non/unitary minimal models, and Wess-Zumino-Witten models. In particular, the conjecture holds beyond diagonal cosets.
@@ 1 Introduction and summary
Conjecture. Let $\mathcal{S}_{\text{UV}}$ be the surviving pre-modular fusion category and $\mathcal{S}_{\text{IR}}$ be its image pre-modular fusion subcategory in the infrared diagonal rational conformal field theory. The (simple) rational renormalization group flow stands in bijection with a unique (up to braided equivalence) completely $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}$-anisotropic representative of the Witt equivalence class $\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$
[…]
@@
Remark. If the surviving pre-MFC $\mathcal{S}_{\text{UV}}$ is non-degenerate, i.e., modular, the conjecture claims the representative of $\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$ is completely anisotropic. Equivalently, the connected étale algebra $A\in\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}$ giving it is the maximal one.
Remark. The unique (up to equivalence) existence of completely $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}$-anisotropic representative of $\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$ signals the universality. Since universality is usually defined in terms of Kan extension, it would be interesting to see the explicit relation between the conjecture and the axiom [14].
@@
Relatedly, when a rational RG flow admits RG domain wall/interface or boundary condition, the conjecture is trivially true by the bijection. Nontrivial statement of the conjecture is that all rational RG flows stand in bijection with completely $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}$-anisotropic representative of $\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}]$.
@@
  • Second, the largest Frobenius-Perron dimension $\text{FPdim}_{\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}}(A)$ of a connected étale algebra $A$ is energetically favored. To see this, we first recall the largest Frobenius-Perron dimension minimizes the Frobenius-Perron dimension of the category $(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})_{A}^{0}$ of dyslectic (right) $A$-modules. More precisely, in order to use the formula (A.7), we first remove the transparent simple objects. The resulting pre-MFC $\mathcal{B}$ no longer have nontrivial transparent simple objects (mathematically, this is denoted $A\cap(\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}})^{\prime}=1$), and it is modular. Since the connected étale algebra does not contain nontrivial transparent simple object, $A$ remains intact. For the MFC $\mathcal{B}$, we can apply the formula to get the smallest Frobenius-Perron dimension $\text{FPdim}(\mathcal{B})$. In pseudo-unitary fusion categories, the Frobenius-Perron dimension equals categorical dimension $\text{FPdim}(\mathcal{B})=D(\mathcal{B})^{2}$. An MFC with categorical dimension $D$ is known to contribute entropy [38, 39] by $\ln D$. At temperature $T$ (defined as the length $1/T$ of the compactified Euclidean time), an MFC with categorical dimension $D$ thus contributes free energy by
@@
Rational RG flow Preserved $\mathcal{S}_{\text{UV}}$ $\mathcal{S}:=\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}$ $\mathcal{S}^{\prime}$ Representative of $\mathcal{S}]$ Section $M(5,4)\to M(4,3)$ Ising $\text{Ising}\boxtimes\text{Ising}$ $\text{Vect}_{\mathbb{C}}$ ToricCode 2.1.1 $M(6,5)\to M(5,4)$ $\text{Fib}\boxtimes\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{1}$ $\text{Fib}\boxtimes\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{1}\boxtimes\text{Fib}\boxtimes\text{sVec}$ $\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{1}\boxtimes\text{sVec}$ $\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{1}\boxtimes\text{sVec}$ 2.1.2 $M(7,6)\to M(6,5)$ $su(2)_{4}$ $su(2)_{4}\boxtimes su(2)_{4}$ $\text{Vect}_{\mathbb{C}}$ $\text{Vect}_{\mathbb{C}}$ 2.1.3 $M(8,7)\to M(7,6)$ $\text{sVec}\boxtimes psu(2)_{5}$ $\text{sVec}\boxtimes psu(2)_{5}\boxtimes\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{1}\boxtimes psu(2)_{5}$ $\text{sVec}\boxtimes\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{1}$ $\text{sVec}\boxtimes\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{1}$ 2.1.4 $M(3,5)\to M(2,5)$ Fib $\text{Fib}\boxtimes\text{Fib}$ $\text{Vect}_{\mathbb{C}}$ $\text{Vect}_{\mathbb{C}}$ 2.2.1 $M(15,4)\to M(9,4)$ Ising $\text{Ising}\boxtimes\text{Ising}$ $\text{Vect}_{\mathbb{C}}$ ToricCode 2.2.2 $\widehat{\mathfrak{su}}(3)_{2}\to\widehat{\mathfrak{su}}(3)_{1}$ $\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}$ $\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}\boxtimes\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}$ $\text{Vect}_{\mathbb{C}}$ $\text{Vect}_{\mathbb{C}}$ 2.3.1 $(\widehat{\mathfrak{e_{7}}})_{3}\to(\widehat{\mathfrak{e_{7}}})_{1}$ $\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{-1}$ $\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{-1}\boxtimes\text{Vec}_{\mathbb{Z}/2\mathbb{Z}}^{-1}$ $\text{Vect}_{\mathbb{C}}$ $\text{Vect}_{\mathbb{C}}$ 2.3.2 $(A_{10},E_{6})\to M(4,3)$ Ising $\text{Ising}\boxtimes\text{Ising}$ $\text{Vect}_{\mathbb{C}}$ $\text{Vect}_{\mathbb{C}}$ 2.4 .
@@ 2.1 Unitary discrete series
(Recall $\mathcal{S}:=\mathcal{S}_{\text{UV}}\boxtimes\mathcal{S}_{\text{IR}}$.) Since our connected étale algebra obeys $A\cap\mathcal{S}^{\prime}=1$, we need to find connected étale algebras in $\text{Fib}\boxtimes\text{Fib}$. The classification was done in [44, 33], and the maximal connected étale algebra is given by
@@
Indeed, in view of anyon condensation [45, 46, 47], one finds $A\in\mathcal{S}$ can be condensed. [^5] The simple objects of $\mathcal{S}_{A}$ are given by $m_{1}\cong A,\quad m_{2}\cong x\oplus X\oplus wW,\quad$ $m_{3}\cong yY\oplus zZ,\quad m_{4}\cong yZ\oplus zY,$ $m_{5}\cong y\oplus zX\oplus yW\oplus zW,\quad m_{6}\cong Y\oplus xZ\oplus wY\oplus wZ,\quad$ $m_{7}\cong z\oplus yX\oplus yW\oplus zW,\quad m_{8}\cong Z\oplus xY\oplus wY\oplus wZ,$ $m_{9}\cong w\oplus W\oplus wX\oplus xW\oplus wW,\quad$ $m_{10}\cong yY\oplus zZ\oplus yZ\oplus zY.$ The first four simple objects are deconfined, and the other six are confined. The deconfined simple objects have conformal dimensions $0,0,\frac{1}{2},0$ mod 1. We also found a non-negative integer matrix representation (NIM-rep), but we omit them. The result tells us
@@
$(\mathcal{S}_{A}^{0})_{A^{\prime}}^{0}]=[\mathcal{S}_{A}^{0}]=[\mathcal{S}].$
Therefore, $(\mathcal{S}_{A}^{0})_{A^{\prime}}^{0}]$ gives a unique (up to braided equivalence) completely anisotropic representative of the Witt equivalence class $\mathcal{S}]$. This example further supports our conjecture. [^7] In order to make this example mathematically rigorous, we still need to prove $A\cong 1\oplus xX\oplus wW$ is a (connected) étale algebra. Since $F$-symbols of $su(2)_{4}$ seem unknown, we could not prove this. However, mathematicians would be able to prove this possibly with some codes.
@@ 2.3 Wess-Zumino-Witten model
where $g$ is the dual Coxeter number. The primary operators (and hence symmetry objects) are labeled by affine weights $^{\hat{\mu}}$ in the affine dominant weights
@@ 2.4 $E$-type minimal model $(A_{10},E_{6})$
The $\mathbb{Z}/2\mathbb{Z}$ simple object in Ising always has conformal dimension $\frac{1}{2}$ mod 1. Thus, it trivially satisfies the half-integer condition. Let us look at the non-invertible simple object. It can have conformal dimensions $\frac{2n+1}{16}$ with $n=0,1,\dots,7$. In UV, it has $h_{y}=\frac{31}{16}$. For its image $Y=F(y)$ to satisfy the half-integer condition, the IR conformal dimension should be either $\frac{1}{16}$ or $\frac{9}{16}$. They give MFCs with additive central charges $\frac{1}{2},-\frac{7}{2}$ mod 8, respectively. For the latter to be unitary, the smallest additive central charge is $\frac{9}{2}$, which is larger than UV. This rules out the latter possibility as follows. All CFTs with central charge $0<c<1$ have been classified [24], and they are minimal models. None of minimal models has Ising with $h=\frac{9}{16}$ for the non-invertible simple object, and central charge smaller than $c_{\text{UV}}=\frac{21}{22}$. Therefore, the only possibility is Ising with $h=\frac{1}{16}$ for the non-invertible simple object. The only possibility is the critical Ising model $M(4,3)$.
In fact, we can do a little more; give conjectures on signs of the relevant coupling with the Cardy’s method [72]. With the Lagrangian coupling $\delta S\ni\lambda_{1/11}\int O_{1/11}$, we try to minimize the energy
$E_{a}=\frac{\pi\cdot 21/22}{24(2\tau_{1})^{2}}-\frac{S_{ja}}{S_{1a}}\frac{\lambda_{1/11}}{(2\tau_{a})^{1/11}}$
for a Cardy state $|a\rangle$ with variational parameter $\tau_{a}$. Here, $j$ is the index corresponding to the relevant operator $O_{1/11}$. The ratio of the modular $S$-matrices take the largest value ($\approx 3.228$) for Cardy states with conformal dimensions $0,\frac{31}{16},\frac{7}{2}$, and the smallest value ($\approx-1.398$) for those with $\frac{1}{11},\frac{133}{176},\frac{7}{22}$. Therefore, for positive Lagrangian coupling, the first three Cardy states give (approximate) vacuum, and for negative Lagrangian coupling, the second three Cardy states give threefold degenerate vacua. Since the first case contains the identity, this signals the massless flow. For the second case, the ground state degeneracy (GSD) three is consistent with mathematical results [31, 32, 35]. To summarize, we arrive the conjecture
$(A_{10},E_{6})+\lambda_{1/11}O_{1/11}=\begin{cases}M(4,3)&(\lambda_{1/11}>0),\\ \text{TQFT w/ GSD}=3&(\lambda_{1/11}<0).\end{cases}$ (2.14)
@@ A.1 Category
$\mathcal{B}^{\prime}:=\{b_{i}\in\mathcal{B}|\forall b_{j}\in\mathcal{B},\ c_{b_{j},b_{i}}\circ c_{b_{i},b_{j}}\cong id_{b_{i}\otimes b_{j}}\}.$ (A.4)
A BFC is called symmetric if all simple objects are transparent. On the other extreme, a BFC is called non-degenerate if the identity object is the only transparent object, $\mathcal{B}^{\prime}=\{1\}$. A non-degenerate BFC $\mathcal{B}\neq\text{Vect}_{\mathbb{C}}$ is called prime if there is no proper non-degenerate braided fusion subcategory other than $\text{Vect}_{\mathbb{C}}$. A generic non-degenerate BFC $\mathcal{B}\neq\text{Vect}_{\mathbb{C}}$ is a Deligne tensor product of prime non-degenerate braided fusion subcategories [26]
@@ A.3 Algebra
Let $\mathcal{B}_{1,2}$ be non-degenerate BFCs. They are called Witt equivalent if there exist fusion categories $\mathcal{C}_{1,2}$ such that $\mathcal{B}_{1}\boxtimes\mathcal{Z}(\mathcal{C}_{1})\simeq\mathcal{B}_{2}\boxtimes\mathcal{Z}(\mathcal{C}_{2})$ is a braided equivalence [26]. This introduces equivalence classes to non-degenerate BFCs. The Witt equivalence class of $\mathcal{B}$ is denoted by $\mathcal{B}$]. For a non-degenerate BFC $\mathcal{B}$ and a connected étale algebra $A\in\mathcal{B}$, we have [26]
$\mathcal{B}]=[\mathcal{B}_{A}^{0}].$ (A.8)
Let $\mathcal{B}_{1,2}$ be non-degenerate BFCs over a symmetric fusion category $\mathcal{E}$ such that the corresponding braided monoidal functors $\mathcal{E}\to\mathcal{B}_{i}\ (i=1,2)$ are fully faithful. The $\mathcal{B}_{1,2}$ are called Witt equivalent if there exist fusion categories $\mathcal{C}_{1,2}$ over $\mathcal{E}$ such that $\mathcal{B}_{1}\boxtimes_{\mathcal{E}}\mathcal{Z}(\mathcal{C}_{1},\mathcal{E})\simeq\mathcal{B}_{2}\boxtimes_{\mathcal{E}}\mathcal{Z}(\mathcal{C}_{2},\mathcal{E})$ is a braided equivalence. It introduces an equivalence relation to non-degenerate BFCs over $\mathcal{E}$. The Witt equivalence class of $\mathcal{B}$ is denoted $\mathcal{B}$]. The set of Witt equivalence classes of non-degenerate BFCs over $\mathcal{E}$ is denoted by $\mathcal{W}(\mathcal{E})$. For a non-degenerate BFC $\mathcal{B}$ over $\mathcal{E}$ and a connected étale algebra such that $A\cap\mathcal{E}=1$, i.e., $A$ does not contain simple objects in $\mathcal{E}$ except the identity object, we have [27]
$\mathcal{B}]=[\mathcal{B}_{A}^{0}].$ (A.9)